# The environment of TeV halo progenitors

### v1

@@ 4.2 Model predictions
Figure 5: Probability for a selection of 30 known pulsars within $2\leavevmode\nobreak\ \mathrm{kpc}$ to be found inside their parent environments, as a function of their characteristic ages. This is based on a Galactic population model made of a proper mix of isolated and cluster massive-star progenitors (see text), and a kick velocity distribution from Igoshev (2020). Dashed and dotted lines mark the $80\%$ and $20\%$ probabilities, respectively.
@@
Figure 6: Probability for a selection of 22 known pulsars within $2\leavevmode\nobreak\ \mathrm{kpc}$ to be found inside their parent environments. Pulsars in that sample have proper motion information $v_{\mathrm{t}}$, which enters as a further constraint on their actual situation (see text). Colored symbols with black edges indicate the probability for $v_{\mathrm{kick}}=v_{\mathrm{t}}$. Boxes indicate the 68% containment interval over possible kick velocities $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$, while whiskers extending above and below the boxes show the 95% containment interval. The horizontal segment inside each box is the median probability.
@@
Before discussing the predictions of our model for the selected known pulsars, we recall that the characteristic age of a pulsar should actually be understood as an upper limit to the true pulsar age. In cases where a more accurate estimate of the age of the pulsar can be obtained by other methods, it appears that the characteristic age can markedly overestimate the true age, by up to one order of magnitude (Suzuki et al. 2021, see also the cases of PSR J0908-4913 and PSR J0538+2817 in Sect. 4.3). In the absence of a reliable estimates for the age of the pulsars in our sample, we use in the following discussion their characteristic age but the caveat above should be kept in mind.
@@
From the age of selected pulsars without velocity information, we can plot Fig. 5, based on the combined WBB+SB case for a Galactic population. There, 8 (8) out of 30 pulsars in that sub-sample have a probability higher than $80\%$ (below $20\%$) of being found inside their parent environment. Such an analysis can be useful to prioritize targets in searches for pulsar halos, depending on the confinement scenario to be tested. The two halo candidates in this sample (J0631+1036, and B0540+23) and the established TeV halo (J0622+3749) have probabilities $\gtrsim 40$% of still residing in their parent environment. In the case of J0631+1036, the probability reaches $\gtrsim 95$%, which makes it a particularly good target to test the idea that the parent environment is providing the conditions for pair confinement.
@@
The sample of pulsars with known transverse velocities is smaller but allows us to extract more information since transverse velocity is a crucial parameters in our model. We still consider for this analysis the combined WBB+SB case for a Galactic population. First, we recall that transverse velocity $v_{\mathrm{t}}$ is the projection on the sky of the kick velocity $v_{\mathrm{kick}}$ of the pulsar such that $v_{\mathrm{kick}}=v_{\mathrm{t}}/\cos(\theta)$ with $\theta$ the angle between the two velocities. A measured $v_{\mathrm{t}}$ therefore results from any value $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$, each implying a different kinematics of the pulsar with respect to the parent environment. The probability of each $v_{\mathrm{kick}}$ is given by the distribution of natal kick velocities from Igoshev (2020), plus a weighing $\propto\cos(\theta)$ to account for the fact that a uniform distribution of kick orientations in space statistically favours velocity vectors close to the plane of the sky (the allowed solid angle is larger for a given $d\theta$ element).
@@
For each pulsar with given characteristic age and transverse velocity, we randomly sample a high number of possible values of $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$ (based on the distributions of kick velocities and orientations mentioned above), and we computed for each $v_{\mathrm{kick}}$ the probability for the pulsar to be located in its parent environment. We can then define ranges for the most frequently encountered probabilities together with confidence intervals. The result is displayed in Fig. 6 in the form of a box plot. For each object in our sub-sample, the following graphical elements are provided: colored symbols with black edges indicate the probability for the smallest possible velocity $v_{\mathrm{kick}}=v_{\mathrm{t}}$; boxes indicate the 68% containment interval, while whiskers extending above and below the boxes show the 95% containment interval; last, the horizontal segment inside each box is the median probability.
@@
Since $v_{\mathrm{t}}$ is the smallest possible kick velocity, and sometimes much smaller than the actual $v_{\mathrm{kick}}$, using it in our model results in a high probability of still being inside the parent environment for most objects, even for quite old pulsars. Applying the above calculation corrects for this bias and avoids overestimating this probability. As a result, and in agreement with previous discussions, pulsars younger than $\tau_{\mathrm{age}}\lesssim 200\leavevmode\nobreak\ \mathrm{kyr}$ still have a high probability of being inside their parent environment while there is a transition from in to out between $200\leavevmode\nobreak\ \mathrm{kyr}\lesssim\tau_{\mathrm{age}}\lesssim 400\leavevmode\nobreak\ \mathrm{kyr}$.
Table 2: Properties of the known pulsars selected for comparison to our model. The probability in the second column is the probability of being found inside the parent environment for pulsars without proper motion information, and the median probability for pulsars with proper motion information, in both cases for the combined WBB+SB scenario for a Galactic population. In the last column, we provide the identification as TeV halo or halo candidate, and as PWN for a normal and fast-moving pulsar (PWN and SPWN, respectively). References: ATNF catalogue (Manchester et al. 2005) except for J0622+3749 (Pletsch et al. 2012), J0633+0632 (Danilenko et al. 2019) and the following parameters,a Khabibullin et al. (2024b), b Van Etten et al. (2012), c Johnston & Lower (2021).
[…]
J0538+2817 99.98 37.6a 360 $4.90\times 10^{34}$ 1.3 PWN
J1809-2332 99.98 67.6b 110 $4.30\times 10^{35}$ 0.88 PWN
B2334+61 99.98 40.6 50 $6.30\times 10^{34}$ 0.7 -
J1044-5737 98.17 40.3 - $8.00\times 10^{35}$ 1.895 -
J0940-5428 97.88 42.2 - $1.90\times 10^{36}$ 0.377 -
J0631+1036 97.41 44.4 - $2.70\times 10^{34}$ 2.104 Candidate Halo
J0554+3107 96.01 51.7 - $5.60\times 10^{34}$ 2 -
J1429-5911 94.12 60.2 - $7.70\times 10^{35}$ 1.96 -
J0633+0632 93.02 59.2 340 $1.20\times 10^{35}$ 1.355 Candidate Halo
J1459-6053 92.94 64.7 - $9.10\times 10^{35}$ 1.84 -
J1954+2836 91.59 69.4 - $1.00\times 10^{36}$ 1.96 -
B0656+14 88.81 111 60 $3.80\times 10^{34}$ 0.286 TeV Halo
B1916+14 85.74 88.1 - $5.10\times 10^{33}$ 1.304 -
J1028-5819 84.96 90 - $8.30\times 10^{35}$ 1.423 -
J0248+6021 84.13 62.4 640 $2.10\times 10^{35}$ 2 TeV Halo
B1727-47 80.03 80.4 500 $1.10\times 10^{34}$ 0.7 -
J1732-3131 78.58 111 - $1.50\times 10^{35}$ 0.641 -
J1740+1000 77.45 114 - $2.30\times 10^{35}$ 1.227 PWN
J2032+4127 62.54 201 20 $1.50\times 10^{35}$ 1.33 -
B0740-28 60.01 157 280 $1.40\times 10^{35}$ 2 SPWN
J0954-5430 59.72 171 - $1.60\times 10^{34}$ 0.433 -
J1301-6310 55.83 186 - $7.60\times 10^{33}$ 1.458 -
B1822-09 53.53 233 20 $4.50\times 10^{33}$ 0.3 -
B0906-49 51.24 112 670 $4.90\times 10^{35}$ 3c SPWN
J0622+3749 50.99 208 - $2.70\times 10^{34}$ 1.6 TeV Halo
J0905-5127 47.99 221 - $2.40\times 10^{34}$ 1.33 -
B0540+23 42.42 253 - $4.10\times 10^{34}$ 1.565 Candidate Halo
J1549-4848 32.90 324 - $2.30\times 10^{34}$ 1.308 -
B0922-52 31.82 335 - $3.40\times 10^{33}$ 0.513 -
B0114+58 29.86 275 154 $2.20\times 10^{35}$ 1.768 -
J1846+0919 29.64 360 - $3.40\times 10^{34}$ 1.53 -
B0941-56 29.04 323 30 $3.00\times 10^{33}$ 0.406 -
J0426+4933 28.77 371 - $2.00\times 10^{33}$ 1.795 -
B0959-54 24.26 441 - $6.90\times 10^{32}$ 0.3 -
B0940-55 22.75 464 - $3.10\times 10^{33}$ 0.3 -
J1849-0317 22.68 481 - $2.90\times 10^{33}$ 1.212 -
J2030+4415 19.87 555 - $2.20\times 10^{34}$ 0.72 -
J2017-2737 19.11 583 - $2.10\times 10^{34}$ 1.549 -
J0633+1746 14.21 342 150 $3.20\times 10^{34}$ 0.19 TeV Halo, SPWN
J1957+5033 13.30 838 - $5.30\times 10^{33}$ 1.365 -
J0835-3707 12.61 877 - $2.40\times 10^{33}$ 0.555 -
J0212+5222 12.56 904 - $4.90\times 10^{33}$ 1.558 -
J1152-6012 12.39 893 - $4.90\times 10^{33}$ 1.367 -
J1848+0647 12.18 916 - $2.70\times 10^{33}$ 1.128 -
J1530-5327 11.74 944 - $8.50\times 10^{33}$ 1.123 -
J1741-2054 7.27 386 160 $9.50\times 10^{33}$ 0.3 SPWN
B0727-18 3.39 426 180 $5.60\times 10^{33}$ 2 -
B1055-52 2.85 535 20 $3.00\times 10^{34}$ 0.093 -
B1742-30 2.07 546 50 $8.50\times 10^{33}$ 0.2 -
B0355+54 1.30 564 60 $5.40\times 10^{34}$ 1 SPWN
B0919+06 0.06 497 460 $6.80\times 10^{33}$ 1.1 -
J0357+3205 $\lesssim$0.01 540 650 $5.90\times 10^{33}$ 0.835 SPWN
B1133-55 $\lesssim$0.01 702 320 $6.70\times 10^{33}$ 1.518 -
@@ 4.3 Discussion of individual pulsars
We provide here a more detailed discussion of a number of specific pulsars and their most likely environment in view of our model’s predictions. We focussed on the pulsars in our sample that are associated with a PWN, because those were the targets of in-depth studies which provided more information about the context of the system (for instance the association with an SNR, the evolutionary state or system age, or the system viewing geometry). We cross-matched our selection of pulsars with the list of PWNe provided in Olmi (2023), keeping the distinction made between normal and fast-moving pulsars (translated into the PWN and SPWN identifiers in Table 2). Eventually, our selection includes nine pulsars with PWNe.
We aimed at classifying these nine pulsars into the following two categories : those that can be located with respect to their putative parent remnant, and those that exhibit typical signatures of a propagation in the ISM (in principle, a pulsar can belong to the two categories, but this did not happen here).
@@
The first category is well-defined, while the second category is less straightforward. The association of a pulsar with a bow-shock pulsar wind nebula (BSPWN) could in principle be indicative that it is travelling in the ISM. Indeed, prototypical BSPWNe are expected to develop when pulsars propagate in a cold and mostly neutral ISM with a relatively low sound speed, and they are characterised by a double shock structure and optical line emissions generated through collisional excitation and charge exchange (Olmi 2023). Yet, the identification to a textbook BSPWN condition is far from simple. There are many circumstances that can give rise to a BSPWN appearance in X-rays or radio (high pulsar proper motion and/or recent interaction with the reverse shock of the remnant), while the typical signatures of a bow shock are not easily detectable or interpretable (complex geometry, low-surface-brightness extended emission, lack of optical emission lines owing to upstream ionization by the pulsar itself). Each case requires a dedicated investigation.
@@
Six pulsars in our sample are listed in Kargaltsev et al. (2017) or Olmi (2023) as fast-moving pulsars with morphological features that are reminiscent of BSPWNe, and three are associated with more classical PWNe. Out of these, three display strong evidence for a connection with an SNR (J1809-2332, J0538+2817, and J0908-4913), to which we added B0656+14 (for reasons detailed below), and two seems to be traveling in the ISM (J0742-2822 and J1741-2054). We discard J0357+3205, J1740+1000, and J0358+5413 because none of the available information on these pulsars can be used in the present context (and the first two are high-latitude objects that may be not representative of the pulsar population). Finally, we provide a specific discussion of J0633+1746, the Geminga pulsar and prototypical halo.
4.3.1 Pulsars likely associated to a SNR
@@
PSR J1809-2332 (powering the Taz PWN): The pulsar is associated with SNR G7.5-1.7, detected as a partial radio shell (Roberts & Brogan 2008). It lies in projection within the remnant, at about half a radius from the center. The proper motion measured from a decade of X-ray observations points back towards the centre of the SNR and suggests a kinematic age of $48\leavevmode\nobreak\ \mathrm{kyr}$ (Van Etten et al. 2012, about 30% less than the characteristic spin-down time). The picture is fully consistent with the prediction of our model, which places the pulsar inside its parent environment with a high probability of $80-100$%.
@@
J0538+2817: The pulsar is associated with SNR Sim 147 and lies well within it in projection, even accounting for the estimated total kick velocity (see Khabibullin et al. 2024a, and references therein). The angular displacement from the geometrical centre of the remnant suggests a kinematic age of $38\leavevmode\nobreak\ \mathrm{kyr}$, much smaller than the characteristic spin-down time of $618\leavevmode\nobreak\ \mathrm{kyr}$. We therefore adopted the former value in lieu of the latter. Interestingly, the properties of the SNR suggest that the explosion of the progenitor star took place in a wind-blown cavity (Khabibullin et al. 2024b), similar to our WBB case or to our SB case for a low-mass cluster. For such a small age, our model unambiguously places the pulsar inside its parent environment with a high probability.
@@
PSR J0908-4913 (also named B0906-49): Johnston & Lower (2021) updated the distance of the pulsar from $1\leavevmode\nobreak\ \mathrm{kpc}$ in the ATNF catalogue to $3\leavevmode\nobreak\ \mathrm{kpc}$, which increases the transverse velocity from $240\leavevmode\nobreak\ \mathrm{km/s}$ to $670\leavevmode\nobreak\ \mathrm{km/s}$. An SNR is associated to PSR J0908-4913 and the pulsar lies in projection on its edge, on the verge of escaping it. Interestingly, our model predicts a $\sim 50$% probability of being inside its parent environment. Yet, there are caveats in this comparison. First, the association with the SNR suggests a kinematical age of $12.7\leavevmode\nobreak\ \mathrm{kyr}$, significantly lower than the characteristic age of $110\leavevmode\nobreak\ \mathrm{kyr}$ of the pulsar (that we used in Fig. 6). Reducing the age of the system to such a lower value would increase the probability of being inside the parent environment up to almost 100%. On the other hand, the physical size of the putative SNR is $10\leavevmode\nobreak\ \mathrm{pc}$ at a distance of $3\leavevmode\nobreak\ \mathrm{kpc}$, which is smaller than the values we obtain at around $10\leavevmode\nobreak\ \mathrm{kyr}$, even in the WBB case. This would tend to pull the probability to lower values. A dedicated modelling of this object with different parameters (lower explosion energy or wind mass loss rate for instance) would be required to get a more solid conclusion. But overall the model predicts probabilities of residing in the parent environment in the range $50-100$%, which is consistent with observations of a pulsar close to the forward shock of its SNR.
@@
PSR B0656+14: this pulsar lies in projection inside the Monogem ring, which has been suggested to be its parent SNR (Knies et al. 2018). With its characteristic age of $110\leavevmode\nobreak\ \mathrm{kyr}$, PSR B0656+14 has a median probability of $90\%$ of still being in its parent environment in our model, which would thus be consistent with that picture. The fact that a TeV halo was detected around B0656+14 strongly suggests that the confinement of particles on the scale of several tens of parsecs has to do with the conditions inside the remnant or bubble. We notice that the radius of the SNR and ISM density are respectively $r=87\leavevmode\nobreak\ \mathrm{pc}$ and $n_{\mathrm{H}}=4\times 10^{-3}$ H cm-3, for a distance of $300\leavevmode\nobreak\ \mathrm{pc}$ and using a Sedov-Taylor evolutionary solution. Such low densities can be expected from SB cavities, and it is therefore possible that the progenitor of B0656+14 exploded in a SB environment, which enabled the SNR to grow to such a large size without being stopped by the SB shell that lies at $\gtrsim 100$ pc (see Fig. 2).

### v2

@@ 4.2 Model predictions
Figure 5: Probability for a selection of 30 known pulsars within $2\penalty 10000\ \mathrm{kpc}$ to be found inside their parent environments, as a function of their characteristic ages. This is based on a Galactic population model made of a proper mix of isolated and cluster massive-star progenitors (see text), and a kick velocity distribution from Igoshev (2020). Dashed and dotted lines mark the $80\%$ and $20\%$ probabilities, respectively.
@@
Figure 6: Probability for a selection of $2$ known pulsars within $2\penalty 10000\ \mathrm{kpc}$ to be found inside their parent environments. Pulsars in this sample have proper motion information $v_{\mathrm{t}}$, which enters as a further constraint on their actual situation (see text). Coloured symbols with black edges indicate the probability for $v_{\mathrm{kick}}=v_{\mathrm{t}}$. Boxes indicate the 68% containment interval over possible kick velocities $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$, while whiskers extending above and below the boxes show the 95% containment interval. The horizontal segment inside each box is the median probability.
@@
Before discussing the predictions of our model for the selected known pulsars, we recall that the characteristic age of a pulsar should actually be understood as an upper limit to the true pulsar age. In cases where a more accurate estimate of the age of the pulsar can be obtained by other methods, it appears that the characteristic age can markedly overestimate the true age, by up to one order of magnitude (Suzuki et al. 2021, see also the cases of PSR J0908-4913 and PSR J0538+2817 in Sect. 4.3). In the absence of a reliable estimate for the age of the pulsars in our sample, we use their characteristic age in the following discussion; nonetheless, the caveat above should be kept in mind.
@@
From the age of selected pulsars without velocity information, we can plot Fig. 5 based on the combined WBB and SB case for a Galactic population. There, 13 (one) out of 30 pulsars in that sub-sample have a probability higher than $80\%$ (below $20\%$) of being found inside their parent environment. Such an analysis can be useful to prioritise targets in searches for pulsar haloes, depending on the confinement scenario to be tested. The two halo candidates in this sample (J0631+1036, and B0540+23) and the established TeV halo (J0622+3749) have probabilities $\gtrsim 65$% of still residing in their parent environment. In the case of J0631+1036, the probability reaches $\gtrsim 95$%, making it a particularly good target for testing whether the parent environment provides the conditions for pair confinement.
@@
The sample of pulsars with known transverse velocities is smaller but allows us to extract more information since transverse velocity is a crucial parameters in our model. We still consider for this analysis the combined WBB and SB case for a Galactic population. First, we recall that transverse velocity $v_{\mathrm{t}}$ is the projection on the sky of the kick velocity $v_{\mathrm{kick}}$ of the pulsar such that $v_{\mathrm{kick}}=v_{\mathrm{t}}/\cos(\theta)$ with $\theta$ the angle between the two velocities. A measured $v_{\mathrm{t}}$ therefore results from any value $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$, each implying a different kinematics of the pulsar with respect to the parent environment. The probability of each $v_{\mathrm{kick}}$ is given by the distribution of natal kick velocities from Igoshev (2020), plus a weighing $\propto\cos(\theta)$ to account for the fact that a uniform distribution of kick orientations in space statistically favours velocity vectors close to the plane of the sky (the allowed solid angle is larger for a given $d\theta$ element).
@@
For each pulsar with a given characteristic age and transverse velocity, we randomly sampled a high number of possible values of $v_{\mathrm{kick}}\geq v_{\mathrm{t}}$ (based on the distributions of kick velocities and orientations mentioned above), and we computed for each $v_{\mathrm{kick}}$ the probability for the pulsar to be located in its parent environment. We then defined ranges for the most frequently encountered probabilities together with confidence intervals. The result is displayed in Fig. 6 in the form of a box plot. For each object in our sub-sample, the following graphical elements are provided: coloured symbols with black edges indicate the probability for the smallest possible velocity $v_{\mathrm{kick}}=v_{\mathrm{t}}$; boxes indicate the 68% containment interval, while whiskers extending above and below the boxes show the 95% containment interval; last, the horizontal segment inside each box is the median probability.
@@
Since $v_{\mathrm{t}}$ is the smallest possible kick velocity and is sometimes much smaller than the actual $v_{\mathrm{kick}}$, using it in our model leads to a high probability that most objects, even quite old pulsars, remain inside their parent environment. Applying the above calculation corrects for this bias and avoids overestimating this probability. As a result and in agreement with previous discussions, pulsars younger than $\tau_{\mathrm{age}}\lesssim 300\penalty 10000\ \mathrm{kyr}$ still have a high probability of being inside their parent environment, while there is a transition from inside to outside between $300\penalty 10000\ \mathrm{kyr}\lesssim\tau_{\mathrm{age}}\lesssim 600\penalty 10000\ \mathrm{kyr}$. All pulsars associated with a TeV halo or a halo candidate have high probabilities of still residing inside their parent environment according to our model.
@@ 4.3 Discussion of individual pulsars
We provide here a more detailed discussion of a number of specific pulsars and their most likely environment in view of our model’s predictions. We focussed on the pulsars in our sample associated with a PWN, because those were the targets of in-depth studies, which provide more information about the context of the system (for instance the association with a SNR, the evolutionary state or system age, or the system viewing geometry). We cross-matched our selection of pulsars with the list of PWNe provided in Olmi (2023), keeping the distinction made between normal and fast-moving pulsars (translated into the PWN and SPWN identifiers in Table LABEL:table:pulsar_properties). Our selection includes nine pulsars with PWNe. We aimed at classifying these nine pulsars into the following two categories : those that can be located with respect to their putative parent remnant, and those that exhibit typical signatures of a propagation in the ISM (in principle, a pulsar can belong to the two categories, but this did not happen here).
@@
The first category is well-defined, while the second category is less straightforward. The association of a pulsar with a bow-shock pulsar wind nebula (BSPWN) could in principle indicate that it is travelling in the ISM. Indeed, prototypical BSPWNe are expected to develop when pulsars propagate in a cold and mostly neutral ISM with a relatively low sound speed, and they are characterised by a double shock structure and optical line emissions generated through collisional excitation and charge exchange (Olmi 2023). Yet, the identification to a textbook BSPWN condition is far from simple. There are many circumstances that can give rise to a BSPWN appearance in X-rays or radio (high pulsar proper motion and/or recent interaction with the reverse shock of the remnant), while the typical signatures of a bow shock are not easily detectable or interpretable (complex geometry, low-surface-brightness extended emission, and lack of optical emission lines owing to upstream ionisation by the pulsar itself). Each case requires a dedicated investigation.
@@
Six pulsars in our sample are listed in Kargaltsev et al. (2017) and Olmi (2023) as fast-moving pulsars with morphological features reminiscent of BSPWNe, and three are associated with more classical PWNe. Out of these, three display strong evidence for a connection with a SNR (J1809-2332, J0538+2817, and J0908-4913), to which we added B0656+14 (for reasons detailed below), while two seem to be travelling in the ISM (J0742-2822 and J1741-2054). We discard J0357+3205, J1740+1000, and J0358+5413 since none of the available information on these pulsars can be used in the present context (the first two are high-latitude objects that may be not representative of the pulsar population). Finally, we provide a specific discussion of J0633+1746, the Geminga pulsar and the prototypical halo.
4.3.1 Pulsars likely associated with a SNR
PSR J1809-2332
@@ 4.3.1 Pulsars likely associated with a SNR
The pulsar is associated with SNR G7.5-1.7, detected as a partial radio shell (Roberts and Brogan 2008), and powers the Taz PWN. It lies in projection within the remnant, at about half a radius from the centre. The proper motion measured from a decade of X-ray observations points back towards the centre of the SNR and suggests a kinematic age of $48\penalty 10000\ \mathrm{kyr}$ (Van Etten et al. 2012, about 30% less than the characteristic spin-down time). The picture is fully consistent with the prediction of our model, which places the pulsar inside its parent environment with a high probability of 9510095-100%.
J0538+2817
@@
The pulsar is associated with SNR Sim 147 and lies well within it in projection, even accounting for the estimated total kick velocity (see Khabibullin et al. 2024a, and references therein). The angular displacement from the geometrical centre of the remnant suggests a kinematic age of $38\penalty 10000\ \mathrm{kyr}$, which is much smaller than the characteristic spin-down time of $618\penalty 10000\ \mathrm{kyr}$. We therefore adopted the former value in lieu of the latter. Interestingly, the properties of the SNR suggest that the explosion of the progenitor star took place in a wind-blown cavity (Khabibullin et al. 2024b), similar to our WBB case or to our SB case for a low-mass cluster. For such a small age, our model unambiguously places the pulsar inside its parent environment with a high probability.
PSR J0908-4913/B0906-49
@@
Johnston and Lower (2021) updated the distance of the pulsar from $1\penalty 10000\ \mathrm{kpc}$ in the ATNF catalogue to $3\penalty 10000\ \mathrm{kpc}$, which increases the transverse velocity from $240\penalty 10000\ \mathrm{km/s}$ to $670\penalty 10000\ \mathrm{km/s}$. A SNR is associated with PSR J0908-4913, and the pulsar lies in projection on its edge, on the verge of escaping it. Our model predicts a $\sim 88$% probability of being inside its parent environment. Yet, there are caveats in this comparison. First, the association with the SNR suggests a kinematical age of $12.7\penalty 10000\ \mathrm{kyr}$, significantly lower than the characteristic age of $110\penalty 10000\ \mathrm{kyr}$ of the pulsar (that we used in Fig. 6). Reducing the age of the system to such a lower value would increase the probability of it being inside the parent environment up to almost 100%. On the other hand, the physical size of the putative SNR is $10\penalty 10000\ \mathrm{pc}$ at a distance of $3\penalty 10000\ \mathrm{kpc}$, which is smaller than the values we obtain at around $10\penalty 10000\ \mathrm{kyr}$, even in the WBB case. This would tend to pull the probability to lower values. A dedicated modelling of this object with different parameters (lower explosion energy or wind mass loss rate for instance) would be required to get a more solid conclusion. But overall the model predicts probabilities of residing in the parent environment higher than 88%, which is consistent with observations of a pulsar in the downstream of the forward shock of its SNR.
PSR B0656+14
@@
This pulsar lies in projection inside the Monogem ring, which has been suggested to be its parent SNR (Knies et al. 2018). With its characteristic age of $110\penalty 10000\ \mathrm{kyr}$, PSR B0656+14 has a median probability of $91\%$ of still residing in its parent environment in our model, consistent with this picture. The fact that a TeV halo was detected around B0656+14 strongly suggests that the confinement of particles on the scale of several tens of parsecs has to do with the conditions inside the remnant or bubble. We notice that the radius of the SNR and ISM density are respectively $r=87\penalty 10000\ \mathrm{pc}$ and $n_{\mathrm{H}}=4\times 10^{-3}$ H cm-3, for a distance of $300\penalty 10000\ \mathrm{pc}$ and using a Sedov-Taylor evolutionary solution. Such low densities can be expected from SB cavities, and it is therefore possible that the progenitor of B0656+14 exploded in a SB environment, which enabled the SNR to grow to such a large size without being stopped by the SB shell lying at $\gtrsim 100$ pc (see Fig. 2).